Be careful to keep a lot of significant figures for all of your answers!

2 R 4 R3

R1 = 2 Ω R2 = 6 Ω R3 = 13 Ω R4 = 8 Ω V = 4 V

1) What is the current through the resistor, R1in the above circuit?

I =

2) What single, equivalent resistor could replace all of the resistors in this circuit?

Rtot =

3)

3) What is the current supplied by the battery?

I =

4) What is the current through the resistor, R2?

I =
Answer #1

R2 and R4 are in parallel

1/6+1/8= 14/48

R' = 48/14= 3.428ohm

R' and R1 are in series R"=2+3.428= 5.428ohm

R" AND R3 ARE IN PARALLEL

1/Req= 1/13 +1/5.428

Req= 2.7126ohm

Current supplied by battery = 4÷2.7126= 1.4746A

Current through R1 and R'

V=4V

4/5.428= 0.7369A

V across R2 is 0.7369 ×3.428= 2.526V

Current through R2

2.526/6= 0.421A